# Machine Learning for Network Attacks Classification and Statistical Evaluation of Adversarial Learning Methodologies for Synthetic Data Generation

### 18 Mar 2026 v1

@@ II-A Data Collection
For all the results that will be presented in the next sections we used the under-sampled version of the UMNIDS dataset, a newly created dataset, which was created by processing raw pcap files from CIC-IDS 2017, CIC-IOT 2023, UNSW-NB15 and CIC-DDoS 2017 [34]. UMNIDS includes flow-level and payload-based data with features that capture historical and temporal patterns in network traffic. This dataset includes around $170,000$ observations, $88$ features and $1$ target variable, which implies if a record is normal or if it is an attack in the network along with the type of the attack. Some that are included are Worms, Bots, DDoS, Infiltration and SQL-Injection. The target variable includes a total of $44$ unique values, $43$ of which specify the kind of attack and the one remaining the non-malicious records. One of the main issues with this dataset is the lack of records for some attacks, like Heartbleed and Infiltrations.
@@ II-B Data Pre-Processing
The pre-processing procedure started by omitting the duplicate values, then encoding the label variable, testing different methods like binning for certain numeric variables with a small number of unique values. We performed an in-depth statistical analysis that included the discovering of linear and non-linear relationships, hypothesis testing for existence of normality through QQ-Plots and goodness of fit tests like Kolmogorov-Smirnov and Anderson Darling, as well as regression analysis for some numeric features. We also examined the existence of outliers with IQR, Z-score, DBSCAN and Local-Outlier Factor and due the great number of outliers we preferred to keep them and use as a scaling method the Robust scaler, after we performed an $80/20$ train/test stratified split, with respect to the attack type variable.
@@ III-B Training Adversarial Models
For the training of the adversarial networks we tested a few neural network architectures, we consider a total number of $20$ epochs for each of the above-mentioned algorithms. From the minimax procedure, for illustration purposes we can see below in Fig. 2 the discriminator’s and generator’s loss values for a few batches, for the Vanilla-GAN and WGAN, from which we observe the failure of Vanilla-GAN as the discriminator’s loss converges to zero and the stable training WGAN, along with indications of convergence:

### 01 Apr 2026 v2

@@ II-A Data Collection
For all the results that will be presented in the next sections we used the under-sampled version of the UMNIDS dataset, a newly created dataset, which was created by processing raw pcap files from CIC-IDS 2017, CIC-IOT 2023, UNSW-NB15 and CIC-DDoS 2017 [34]. UMNIDS includes flow-level and payload-based data with features that capture historical and temporal patterns in network traffic. This dataset includes around $170,000$ observations, $88$ features and 11 target variable, which implies if a record is normal or if it is an attack in the network along with the type of the attack. Some that are included are Worms, Bots, DDoS, Infiltration and SQL-Injection. The target variable includes a total of $44$ unique values, $43$ of which specify the kind of attack and the one remaining the non-malicious records. One of the main issues with this dataset is the lack of records for some attacks, like Heartbleed and Infiltrations.
@@ II-B Data Pre-Processing
The pre-processing procedure started by omitting the duplicate values, then encoding the label variable, testing different methods like binning for certain numeric variables with a small number of unique values. We performed an in-depth statistical analysis that included the discovering of linear and non-linear relationships, hypothesis testing for existence of normality through $Q$-Plots and goodness of fit tests like Kolmogorov-Smirnov and Anderson Darling, as well as regression analysis for some numeric features. We also examined the existence of outliers with IQR, Z-score, DBSCAN and Local-Outlier Factor and due the great number of outliers we preferred to keep them and use as a scaling method the Robust scaler, after we performed an $80/20$ train/test stratified split, with respect to the attack type variable.
@@ III-B Training Adversarial Models
For the training of the adversarial networks we tested a few neural network architectures, we consider a total number of 2020 epochs for each of the above-mentioned algorithms. From the minimax procedure, for illustration purposes we can see below in Fig. 2 the discriminator’s and generator’s loss values for a few batches, for the Vanilla-GAN and WGAN, from which we observe the failure of Vanilla-GAN as the discriminator’s loss converges to zero and the stable training WGAN, along with indications of convergence: